# Problem Solved? Information Extraction Design Space for Layout-Rich Documents using LLMs

### v1

@@ Appendix A Appendix
A.1 Prompt Generation Details
@@
Figure 4: Few-Shot Prompt Structure with 1-shot Example
Figure 5: Chain of Thought Prompt Structure with 1-shot Example
A.2 Tailoring Dataset for our Test-Suite
We used the Registration Form with six entity types for this project, as shown in Figure 6.
@@ A.2 Tailoring Dataset for our Test-Suite
The VRDU dataset includes predefined few-shot splits consisting of training, testing, and validation sets. These splits contain 10, 50, 100, and 200 training samples, each with three subsets, as shown in Figure 7. The dataset also includes different levels (Lv1: Single, Lv2: Mixed, and Lv3: Unseen Type) and various template types (Amendment, Dissemination, and Short-Form).
@@
For the models GPT-3.5, GPT-4o, and LLaMA3-70B, we created these few-shot and CoT examples. However, no few-shot examples were used for GPT-4 Vision; we used the same test dataset as for the other three models with basic instructions. For LayoutLMv3, we used the same training and test datasets as the other models but included the entire validation set (300 samples) instead of selecting only five. The training and test datasets remained unchanged across all models to ensure a fair evaluation and consistency.
A.3 Success of Evaluation Techniques
@@
Figure 9: Comparison of error rates in Substring vs. Fuzzy Matching
Based on our intubation, we came up with seven different error categories see Figure 9. From left to right in Error Types 5 OCR error means errors occur because of OCR; for example, handwritten names’ or signatures are sometimes extracted as different names incorrectly (e.g Jim Slattery handwriting extracted as Jim Slatters) GT error represents the case where GT information is not completely true sometimes its missing some words and sometimes it has additional words for example gt value is ’Om Saudi Arabia 1’ but it is Kingdom’ Saudi Arabia in the document. LLM Halucination represents where LLM come up a prediction that is not exist data in the OCR for example gt value is for file date ’1992-04-21’ but prediction is ’1992-04-24’ even though OCR data and document itself does not contain any number 24. Additional info represents that prediction is correct, but there is some additional info’ for example, while the signer’s name is ’Daniel Manatt” LLM predicts it as ’Daniel Manatt Todd’ where Todd exists in OCR and the document has a notary name as a handwritten signature. Wrong info means gt is exist in the OCR and the document but LLM choose another date as prediction for example while file date ’2016-10-08’ exist in OCR LLM return 2016-10-31 as file date which belongs registration date in the document.
@@
Human error is for the errors occurred physically by humans failures for example sometimes documents ar scanned partly/crooked by humans. Finally incomplete predictions represent prediction is include gt but not completely for example sometime gt is ’Japan External Trade Organization Tokyo, Japan’ but prediction is missing only last part which is Tokyo, Japan. Figure 9 represent all these error types and their rates in both Substring and Fuzzy match categories where exact match is 0 but they labeled as 1.
For our evaluation the success of Fuzzy and Substring we decided to label data points as 0 only in the case where wrong info is 1. For other error categories we accept them and label gt as 1 for those. Table 11 represent performance metrics for fuzzy and substring calculated with gt labels explained above with the data points (exact match 0 substring/fuzzy 1)

### v2

@@ Appendix A Appendix
A.1 Discussion & Observations
Effect of Dataset Characteristics on Parameter Selection.
The differences in parameter selection between the VRDU (OCR–Max–Few-shot–0–Cleaned Pred.–Fuzzy Match) and FUNSD (Markdown–Small–Few-shot–0–Cleaned Pred.–Fuzzy Match) datasets stem from their inherent characteristics. VRDU contains long, multi-page documents extracted via OCR, often resulting in noisy and fragmented text, where using the raw OCR output preserves the original structure. In contrast, FUNSD consists of structured, labeled question-answer pairs, making Markdown a suitable input format for key-value pair extraction. These differences shape the parameter selection process, as some models may inherently favor one format over another due to training data characteristics—e.g., GPT-4o may handle OCR data more effectively, while Llama3 shows a preference for Markdown.
Unexpected Impact of Markdown Representation on Model Performance.
Markdown, while useful for structuring data, does not always improve performance when applied to noisy and unstructured inputs like those from the VRDU dataset. We observed that converting OCR outputs to Markdown could introduce inconsistencies, such as misaligned headers, broken paragraphs, or flattened tables, affecting data representation quality. In contrast, retaining the raw OCR output often preserved the original structure. For structured datasets like FUNSD, however, Markdown aligned well with the data format, facilitating key-value pair extraction. After experimenting with the IE pipeline on VRDU using purely text input without bounding box locations from OCR, we achieved an overall F1 score of 0.632 (with OCR and Markdown results presented in Table 2). These findings demonstrate that (1) LLMs can leverage structural information from both OCR and Markdown, and (2) their performance varies depending on the input format. These results suggest that the effectiveness of Markdown may depend on both dataset characteristics and model preferences, as previously discussed.
Insights into the Effects of Few-Shot Examples and Chain-of-Thought Prompting.
Few-shot prompting with zero examples consistently outperforms other configurations due to the model’s ability to generalize without being constrained by specific instances. In VRDU, the variability and noise introduced by OCR, along with the lack of a fixed key-value schema in the FUNSD dataset, make it challenging for the model to generalize from a few specific examples. Few-shot examples often include multiple key-value pairs within a single text that may not appear in other documents, resulting in a mismatch between training and test data. Chain-of-Thought (CoT) prompting, despite its usefulness for reasoning tasks, can overcomplicate straightforward key-value extraction, as guiding the model to think step by step may distract from directly mapping structured content. Moreover, few-shot examples often lack layout information compared to test documents due to the way they were created (see A.3), further contributing to mismatches. Consequently, zero-shot prompting proves more effective by avoiding template-specific biases, maintaining prompt brevity, and enabling better generalization across diverse document structures.
Potential Optimization Bias in Substring and Fuzzy Matching Evaluations.
During manual inspection of ground truth and extracted data, we noticed that exact matching often introduces bias by favoring the original annotation, leading to false errors. For example, ’signer_title’ in the ground truth is ’general manager’, while the input file and extraction have ’general manager, north america’—both correct, yet exact matching would mark the latter as wrong. Similarly, for ’signer_name’, the ground truth has [’Catherine Redman-randell’], but the extraction includes [’A.j. Maltrentin’, ’Catherine Redman-randell’], which is accurate given the context.
To assess the reliability of substring and fuzzy matching, we manually annotated errors, categorizing them as OCR error, ground truth error, LLM hallucination, additional/wrong information, human error, or incomplete prediction. Substring matching reported zero errors, while fuzzy matching had just one error among 1,730 extractions (see Appendix A.4). These findings support the use of fuzzy matching in scenarios where exact matching would unfairly penalize valid variations.
Selection Criteria for OFAT and Brute-Force Configurations in Performance Evaluation.
The best case configurations for each model were determined using a full-factorial (brute-force) exploration, as detailed in Table 8, since the optimal configuration for one model may not generalize well to others. In contrast, for the OFAT configurations, we selected the parameter set that achieved the highest average performance across comparisons—specifically, OCR–Max–Few-shot–0–Cleaned Pred.–Fuzzy Match—as shown in Tables 25.
While it might seem inconsistent to use OFAT for overall results and brute-force for best/worst cases, the rationale is that OFAT aims to identify generally optimal parameters with minimal effort, whereas the brute-force method is necessary to assess each model’s best and worst performance individually for comparison purposes.
A.2 Prompt Generation Details
Figure 4: Few-Shot Prompt Structure with 1-shot Example.
Figure 5: Chain of Thought Prompt Structure with 1-shot Example.
@@ A.2 Prompt Generation Details
A.3 Tailoring Dataset for our Test-Suite
VRDU dataset includes two benchmarks, each including train samples of 10, 50, 100, and 200 documents. The Registration Form with six entity types used for this project, see Figure 6.
@@ A.3 Tailoring Dataset for our Test-Suite
The VRDU dataset includes predefined few-shot splits that consist of train, test, and validation sets. These splits contain 10, 50, 100, and 200 training samples, each with three subsets, as shown in Figure 7. The dataset also includes different levels (Lv1: Single, Lv2: Mixed, and Lv3: Unseen Type) and various template types (Amendment, Dissemination, and Short-Form).
@@
For GPT-3.5, GPT-4o, and LLaMA3-70B, we used few-shot and CoT examples, while GPT-4o-vision used the same test dataset without few-shot examples, following basic instructions. For LayoutLMv3 and ERNIE-Layout, we used the same training and test datasets as the other models but included the entire validation set (300 samples). This consistent setup ensured fair evaluation across all models.
A.4 Success of Evaluation Techniques
@@
Figure 9: Comparison of error rates in Substring vs. Fuzzy Matching.
Based on our intubation, we identified seven error categories (see Figure 9): OCR errors occur when handwriting is misinterpreted (e.g., "Jim Slattery" as "Jim Slatters"). GT errors arise from incomplete or inaccurate ground truth (e.g., "Om Saudi Arabia 1" instead of "Kingdom Saudi Arabia"). LLM hallucinations involve predictions not present in the OCR or document (e.g., predicting "1992-04-24" instead of "1992-04-21"). Additional info errors include correct predictions with extra information (e.g., "Daniel Manatt Todd" instead of "Daniel Manatt"). Wrong info occurs when LLM selects an incorrect value despite the correct one being available (e.g., predicting "2016-10-31" as the file date instead of "2016-10-08").
Human errors result from physical mistakes, such as crooked or incomplete scans. Incomplete predictions occur when the prediction includes part of the ground truth but misses a segment (e.g., "Japan External Trade Organization" instead of "Japan External Trade Organization Tokyo, Japan"). Figure 9 shows these error types and their rates for both Substring and Fuzzy match categories, where exact matches are labeled as 1.
For evaluating the success of Fuzzy and Substring, we labeled data points as 0 only when the error category is "wrong info"; for other error types, we accepted them and labeled the ground truth as 1. Table 10 presents the performance metrics for Fuzzy and Substring calculated based on these ground truth labels, where exact match is 0 and substring/fuzzy is 1.
@@
A.5 Additional Experiments with FUNSD Dataset
To demonstrate the generalizability of our techniques, we applied the same steps to the FUNSD dataset as to VRDU.
Table 11: Performance results for different LLMs on the FUNSD dataset with different input types. Baseline configuration in light blue.
Models Exact Match (F1)
OCR Markdown
GPT-3.5 0.357 0.457(+0.100)
GPT-4o 0.498 0.523(+0.025)
LLaMA3 0.387 0.490(+0.103)
Avg ($\pm$ stdev.) 0.414($\pm$ 0.074) 0.490($\pm$ 0.033)
For the FUNSD dataset, Markdown input consistently outperforms OCR across all models. GPT-3.5 and LLaMA3 show the most significant improvements (+0.100 and +0.103, respectively), while GPT-4o sees a smaller gain (+0.025). On average, Markdown performs better (0.490 vs. 0.414), with lower variability, indicating its effectiveness for structured data in FUNSD (see Table 11).
Table 12: Performance results for different LLMs on the FUNSD dataset with different chunk size categories. Baseline configuration in light blue.
Models Exact Match (F1)
Small($\leq 1024$) Medium($\leq 2048$) Max($\leq 4096$)
GPT-3.5 0.354(-0.003) 0.357 0.352(-0.005)
GPT-4o 0.503(+0.005) 0.498 0.494(-0.004)
LLaMA3 0.390(+0.003) 0.387 0.389(+0.002)
Avg ($\pm$ stdev.) 0.415($\pm$ 0.77) 0.414($\pm$ 0.074) 0.411($\pm$ 0.073)
Small chunk size yields a slight performance advantage on FUNSD with average F1 of 0.415 over medium (0.414) and max (0.411), likely due to its alignment with FUNSD’s short Q&A pairs–preserving structure, reducing context dilution, and focusing models on concise, relevant segments (Table 12).
Table 13: Different LLMs with different prompt types and example numbers. Baseline configuration is highlighted in light blue.
Models Exact Match (F1)
0 1 3 5
few-shot
GPT-3.5 0.357 0.289(-0.068) 0.283(-0.074) 0.316(-0.041)
GPT-4o 0.498 0.467(-0.031) 0.399(-0.099) 0.417(-0.081)
LLaMA3 0.387 0.397(+0.010) 0.398(+0.011) 0.393(+0.006)
Avg ($\pm$ stdev.) 0.414($\pm$ 0.074) 0.384($\pm$ 0.089) 0.360($\pm$ 0.66) 0.375($\pm$ 0.052)
CoT
GPT-3.5 0.327(-0.030) 0.184(-0.173) 0.205(-0.152) 0.218(-0.139)
GPT-4o 0.482(-0.016) 0.396(-0.102) 0.369(-0.129) 0.358(-0.140)
LLaMA3 0.407(+0.020) 0.377(-0.10) 0.360(-0.027) 0.385(-0.002)
Avg ($\pm$ stdev.) 0.404($\pm$ 0.109) 0.290($\pm$ 0.149) 0.287($\pm$ 0.115) 0.288($\pm$ 0.98)
Zero-shot prompting consistently yields the best performance on the FUNSD dataset, with average F1 scores of 0.414 (±0.074) for few-shot prompting and 0.404 (±0.109) for CoT. Few-shot prompting is generally more stable than CoT, especially for GPT-3.5 and GPT-4o. Adding examples usually reduces performance, likely due to noise and limited generalization. LLaMA3 remains more stable with CoT compared to other models, but overall, few-shot prompting is better suited for FUNSD’s short, structured Q&A format (see Table 13).
The results indicate that Cleaned Prediction consistently yields better or comparable F1 scores compared to Initial Prediction, with minor improvements due to effective noise reduction, while Mapped Prediction consistently reduces performance due to overgeneralization and key misalignment. This issue arises because the FUNSD dataset lacks a fixed schema, causing key-value pairs to vary dynamically between documents, leading to mapping errors (see Table 14).
Table 14: Different LLMs with different post-processing strategies. Baseline configuration is highlighted in light blue.
Models Exact Match (F1)
Initial Pred. Mapped Pred. Cleaned Pred.
GPT-3.5 0.357 0.311(-0.046) 0.361(+0.004)
GPT-4o 0.498 0.456(-0.042) 0.496(-0.002)
LLaMA3 0.387 0.348(-0.039) 0.393(+0.006)
Avg ($\pm$ stdev.) 0.414($\pm$ 0.074) 0.371($\pm$ 0.075) 0.416($\pm$ 0.070)
The results indicate that Substring Match and Fuzzy Match consistently outperform Exact Match across all models, with average F1 scores of 0.488 (±0.060) and 0.497 (±0.092) respectively, compared to 0.414 (±0.074) for Exact Match. This improvement is due to the more flexible matching criteria of substring and fuzzy evaluation, which better accommodate minor variations or partial matches in predictions. The performance gain is most pronounced in GPT-4o, highlighting that models producing slightly varied outputs benefit significantly from more tolerant evaluation metrics (see Table 15).
Table 15: Different LLMs with different evaluation techniques. Baseline configuration is highlighted in light blue.
Models Evaluation Metric (F1)
Exact Match Substring Match Fuzzy Match
GPT-3.5 0.357 0.440(+0.083) 0.423(+0.066)
GPT-4o 0.498 0.557(+0.059) 0.601(+0.103)
LLaMA3 0.387 0.469(+0.082) 0.468(+0.081)
Avg ($\pm$ stdev.) 0.414($\pm$ 0.074) 0.488($\pm$ 0.060) 0.497($\pm$ 0.092)
As OFAT, drawing from the underlying 12 experiments, we identified the optimal parameter for each step and each model based on the experimental outcomes (Table 16). Also, we conducted an exhaustive full factorial exploration with 432 configurations to find the best parametrization per LLM (Table 17). Lastly, we find the worst configuration on a per-LLM and per-model basis.
Parameter GPT-3.5 GPT-4o LLaMA3
Input Type Markdown Markdown Markdown
Chunk Size Medium Small Small
Prompt Few-Shot Few-Shot CoT
Example No. 0 0 0
Output Refin. Cleaned Initial Cleaned
Evaluation Substring Fuzzy Substring
Parameter GPT-3.5 GPT-4o LLaMA3
Input Type Markdown Markdown Markdown
Chunk Size Max Medium Small
Prompt Few-shot CoT Few-shot
Example No. 3 0 0
Output Refin. Cleaned Cleaned Cleaned
Evaluation Fuzzy Fuzzy Fuzzy
Figure 10: F1 scores of different LLMs across three configurations (Baseline, OFAT, and Brute-Force). Each bar represents the mean F1 score of a model for the corresponding configuration.The OFAT configuration uses a single global parameter set optimized across all models.
The performance of these different configurations is depicted in Figure 10. The OFAT method achieves an average F1 score of 0.789, marking a substantial gain of $37.5$ points over the baseline configuration (0.414). In comparison, the full factorial exploration yields a slightly higher average F1 score of 0.797–only $0.8$ points more–while evaluating all 432 configurations. Notably, OFAT achieves this performance by exploring just 12 configurations, requiring approximately 2.8% of the computational effort compared to the full factorial search space. Thus, OFAT captures over 99% of the brute-force method’s effectiveness, making it a highly efficient alternative when computational resources are limited. Overall, OFAT continues to offer a compelling balance between performance and efficiency. In comparison, the worst configuration achieves only 0.226 on average, which is approximately 28% of the best configuration.